OpenStudy (anonymous):

PROJECTILE MOTION I don't know how to get started with this. A grasshopper jumps a horizontal distance of 1m from rest, with an initial velocity at a 45 degree angle with respect to the horizontal. Find (a) the initial speed of the grasshopper and (b) the maximum height reached.

OpenStudy (michele_laino):

here we have to apply the formula of the range \(r\), namely: \[r = \frac{{v_0^2}}{g}\sin \left( {2\theta } \right)\]

OpenStudy (michele_laino):



OpenStudy (anonymous):

Where did you derive that from? Which kinematic equation?

OpenStudy (michele_laino):

that formula expresses the range \(r\) of an object, which is launched with a velocity whose magnitude \(v_0\) and angle (with respect to the horizontal line) equal to \(\theta\)

OpenStudy (michele_laino):

it is a standard formula, which can be derived from the theory of falling bodies

OpenStudy (michele_laino):

in our case we have: \(\theta=45\;degrees\)

OpenStudy (michele_laino):

from my formula above, we get: \[{v_0} = \sqrt {\frac{{rg}}{{\sin \left( {2\theta } \right)}}} = \sqrt {\frac{{1 \cdot 9.81}}{{\sin \left( {2 \cdot 45} \right)}}} = ...m/\sec \]

OpenStudy (anonymous):

I don't seem to understand that. We're only given three general equations.